OpenStudy (anonymous):

In order to dunk a basketball, Jim must jump to a height of 1.2 meters. He can jump with an initial velocity of 5.1m/sec. Can he dunk? (quadratic solution?)

OpenStudy (anonymous):

We know u=5.1, s=? a=9.8 and final velocity = 0 Using \[v^2=u^2+2as\]

OpenStudy (anonymous):

\[0=5.1^2 + 2 (-9.8)s\]

OpenStudy (anonymous):

So a = -9.8 sorry

OpenStudy (anonymous):

Solve for s (displacement) if >1.2 he can dunk :) that ok?

OpenStudy (anonymous):

I'm supposed to use the quadratic formula to solve this one, specifically the discriminant.

OpenStudy (anonymous):

Um

OpenStudy (anonymous):

Yeah, that's what I said!

OpenStudy (amistre64):

the physics quadratic for this is:\[h=-\frac12gt^2+v_ot+h_o\]

OpenStudy (amistre64):

the quadratic formula can tell us when he jumps, and when he lands, the determinant really isnt all the useful to us in determining the vertex of the parabola tho

OpenStudy (amistre64):

i spose when h=1.2 ... that will zero things out in the end

OpenStudy (amistre64):

\[1.2=-4.9t^2+5.1t\] \[4.9t^2-5.1t+1.2=0\] and i see i confused all the d terms, discriminant is the axis of symmetry value

OpenStudy (anonymous):

This question was part of an assignment that started with the outcomes wen the "discriminant" was positive vs. negative in an algebra quadratic formula, so I'm guessing he's driving at my being able to put the information given into that formula. So much for explicit teaching!!

OpenStudy (amistre64):

discriminant is defined as: -b/2a 5.1/2(4.8) = t

OpenStudy (amistre64):

.... my fingers hate me today :)

OpenStudy (anonymous):

Not sure where the 4.8 and 4.9 came from.